OpenStudy (anonymous):

Consider the following unbalanced reaction: C(s) + S8(s) → CS2(l) How many grams of carbon disulfide will be produced if 8.80 g of C and 17.2 g of S8 are allowed to react completely? Report your answer to 3 significant figures and do not include units.

OpenStudy (aaronq):

do you know what units you use in stoichiometric problems?

OpenStudy (anonymous):

in this case g CS2

OpenStudy (anonymous):

MOLEES?

OpenStudy (aaronq):

nope you ALWAYS use moles

OpenStudy (anonymous):

OK

OpenStudy (aaronq):

so to solve this problem: (1) balance the reaction (2) convert grams of reactants to moles of reactants ...moles=mass/molar mass (3) determine limiting reactant (whichever one you have less of will determine how much product you can make) (4) use the coefficients from the balanced equation to find how many moles of product you make (5) convert to grams ...again using n=m/M

OpenStudy (aaronq):

do these and ill help you if you get stuck

OpenStudy (anonymous):

im stuck...im trying to convert the grams of the reactant to moles of reactant but im confused

OpenStudy (aaronq):

okay, so you know how many grams you have .. do you know each reactants molar mass?

OpenStudy (anonymous):

molar mass of C=12.011?

OpenStudy (anonymous):

Molar mass of CS2=76.141

OpenStudy (aaronq):

okay so use the formula n=m/M moles = mass / molar mass find the number of moles ie. divide the mass by the molar mass

OpenStudy (aaronq):

so for carbon m=(8.8g)/(12g/mol) = 0.733 mol

OpenStudy (anonymous):

OK

OpenStudy (preetha):

Aaron, really good work helping this learner!